Question 49
ProbabilityLet P be a probability function on set S, where S = {a1, a2, a3, a4}. Find P{a1} if P{a2} = ⅓, P(a3) = ⅙ and P(a4) = ⅕.
- A.7/10
- B.⅔
- C.⅓
- D.3/10
Show the answer and explanation
Answer: D (3/10)
Explanation
The four probabilities must add up to 1. P(a₁) = 1 − (⅓ + ⅙ + ⅕) = 1 − 21/30 = 9/30 = 3/10.